Status Report

ISS On-Orbit Status 7 Aug 2002

By SpaceRef Editor
August 7, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

On wake-up, the station residents found enthusiastic kudos from MCC-H for their wonderful job handling yesterday’s PAO/educational TV event with the Maryland Science Center.

Crew today continued preparations for next week’s first EVA (8/16), to be conducted by CDR Korzun and FE-1 Whitson in Orlan-M suits from the “Pirs” DC-1 airlock module. [A list of approved and Hydrolab-tested US EVA tools was uplinked by MCC-H for possible use during the two spacewalks.  Korzun and Whitson spent about an hour preparing the Russian EVA tool carrier (KPU), which holds tools with rubber bands and tetherhook bungees.  Korzun and FE-2 Sergei Treschev used another 1.5 hours for careful pre-EVA preparation of the new KROMKA payload that will replace the current KROMKA tablet outside the Service Module (SM).  The activities were recorded with the US DVCAM (digital video camera) for subsequent review by ground specialists, and the crewmembers also tagged up with TsUP (MCC-M) via audio (VHF and S-band).]

Also for the two EVAs, Peggy Whitson prepared and checked out the two important PGTs (pistol-grip tools, i.e., motor-driven bolt drivers) required for outboard work.  [Because the power switch nut on one of the PGTs came loose during the STS-109/Hubble-4 mission, resulting in a loose switch, a special verification procedure was uplinked for Peggy to perform today to ensure this does not happen on the upcoming EVA.  She was requested to do it on all three PGTs as this must be done prior to 9A.]

Whitson also terminated charging of the remaining REBA (rechargeable EVA battery assembly) batteries started yesterday.

CDR Korzun and FE-2 Treschev meanwhile moved the freshly filled portable oxygen repress tank (BNP) from the SM to the Russian DC-1 module and hooked it up inside the “Pirs” airlock.

After the BNP installation, Korzun checked out connections and tested the EVA support panel (POV) in the vestibule tunnel between the SM PkhO Transfer Compartment and the DC-1.

FE-1 Whitson terminated the current microbial air sampling session by retrieving the FMK (formaldehyde monitoring kit) monitors deployed yesterday in the SM and Lab and deactivating the SSAS (solid sorbent air sampler) collectors.

Peggy also completed SUBSA-04, her fourth science run with the SUBSA (solidification using a baffle in sealed ampoules) experiment.  [After first activating the MSG (Microgravity Science Glovebox), she used it to place the specimen in the furnace and then operate the experiment, which studies production of semiconductor materials.  The SUBSA ground team also reacted with excitement when Peggy suggested to make up for the time lost with earlier troubleshooting by assisting with another added run on Saturday,

The crew continued the Renal (Kidney) Stone research program’s current session.  While the CDR’s regimen of urine collection came to an end today, Treschev started on his collection phase plus metabolic log entries, and Whitson began her part in the experiment with dietary intake logging.

All crewmembers also completed their regular weekly psychological NTXN “Interactions” data entry session by completing the prescribed encrypted mood, group and journal questionnaire on the MEC (medical equipment computer).

Later in  the day, Peggy Whitson set up and participated in a ham (amateur) radio pass with Dutch and French youngsters at the Space Camp at Euro Space Center in Transinne, Belgium.  [During the week of 7/7, 40 youngsters from USAF families stationed in Germany attended the Space Camp and wrote 20 questions for the ISS Flight Engineer in English.  Today, Dutch- and French-speaking students asked those questions.  They were provided with a translation of Peggy’s answers, which also will be circulated to the USAF families.]

Troubleshooting on the TVIS
treadmill, currently still off-limits for the crew, is continuing at MCC-H.  Specialists are now more confident that by giving the crew some operational constraints (speed/load/inspections), they may be able to run “as is”.  A final decision will be made by the end of the week.  Until then, the crew is to secure the damaged roller and button up TVIS using a procedure uplinked overnight for the “job jar” task list.  It will take two crewmembers an estimated 1.5 hours.

Daily routine servicing tasks were performed by CDR Korzun (inspection of BRPK-2 water condensate separator, maintenance of SOSh life support system), FE-2 Treschev (IMS delta file preparation), and FE-1 Whitson (status checks of PCG-STES-008 and ADVASC

All crewmembers performed their daily physical exercise on RED (resistive exercise device) and VELO (cycle ergometer with load/strength

Today’s targets for the CEO (crew earth observation) program were Angolan Biomass Burning (this pass offered good pan views of fire activity in Angola to the right of track), High Central Andean Glaciers (optimum lighting and few clouds made this an excellent pass for context views of the small ice fields and glaciers of this region. As ISS crossed the Peruvian coast from the southwest, crew was to look right of track [SE] along the crest of the Andes), Eastern Mediterranean Dust (high pressure and smog conditions should have held over the eastern Med this pass.  Of interest as ISS crossed the Egyptian coast: looking right of track for aerosols from the lower Nile valley and coastal cities of the Middle East), Tigris-Euphrates, Turkey (ground track was to the NW of most of this target area.  Good context views of reservoir and stream patterns should have been visible to the right of track), and Lower Amazon River Basin (on this pass the mouth of the Amazon River was to the left of track.  Crew was to try for context photos using sun glint to enhance water boundaries).
CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 1:05 pm EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp mode).  Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed).  U.S. CDRA CO2 scrubber is Off (leak).  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.

SM Working Compartment:  Pressure (mmHg) — 761, temperature (deg C) — 26.9, ppO2 (mmHg) — 156.1, ppCO2 (mmHg) — 2.6.
SM Transfer Compartment:  Pressure (mmHg) — 761, temperature (deg C) — 20.7.
FGB Cabin:  Pressure (mmHg) — 756, temperature (deg C) — 21.3.
Node:  Pressure (mmHg) — 752.79, temperature (deg C) — 25.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab:  Pressure (mmHg) — 754.82, temperature (deg C) — 25.1, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) — 754.82, temperature (deg C) — 28.6; shell heater temp (deg C) — 27.6, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1:  Shell heater temp (deg C) — 24.3
PMA-2:  Shell heater temp (deg C) — 14.7

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] is failed and in Extended Life mode [=3D a state that preserves mass spectrometer vacuum but produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly) were replaced, but some more work needs to be done).

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack mode; BGA 4B also in Autotrack mode.
SM batteries: Battery #1 is off-line, battery #6 is cycling; all other batteries (6) are in “Partial charge” mode.
FGB batteries:  Battery #5 is off-line, battery #3 is cycling; all other batteries (4) are in “Partial charge” mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:
Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs online.
State vector — US GPS (SIGI string 1)
Attitude — Russian segment
Angular rates — US RGA1 (rate gyro assembly 1)

Communications Systems:
All Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1) and MBS PDGF 4, with Keep Alive power on both strings (based on MBS).
MBS:  Operational on both strings.
RWS (robotics workstations): Cupola is Active, Lab RWS is Off.

ISS Orbit (as of this morning, 6:44 am EDT [ epoch]):
Mean altitude — 395.5 km
Apogee — 407.7 km
Perigee — 383.2 km
Period — 92.5 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0018053
Orbits per 24-hr. day — 15.57
Altitude decrease — 150 m (mean) in last 24 hours
Solar Beta Angle — 24.0 deg (magnitude decreasing)
Revolutions since FGB/Zarya launch (Nov. =9298) — 21202
Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane “sun-fixed” [yaw: ~180 deg, pitch: -5.7 deg., roll: 0 deg]).  Next maneuver to LVLH (“earth-oriented”) on Friday, 8/9.

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.